OpenStudy (anonymous):

Solve equation (2+2i) z^3=-1, where z is complex number.

OpenStudy (shubhamsrg):

solve for ? z ?

OpenStudy (anonymous):

yes

OpenStudy (shubhamsrg):

hmmn,,make thise substitutions: 1+i = sqrt(2)e^(ipi/4) solve for z ..was it helpful ?

OpenStudy (anonymous):

wel... not really

OpenStudy (anonymous):

it only complicates things

OpenStudy (shubhamsrg):

hmmn we have this 2(1+i) z^3 = -1 =>2(_/2)e^(ipi/4) z^3 =-1 =>z= [-1/2(_/2)e^(ipi/4)]^1/3 just simplify.. note that e^(ipi/12) = cos(pi/12) + isin(pi/12)